6. Knowing that F1 = 10 N, d1 = 8 cm, F2 = 8 N and d2 = 5 cm, determine: a) Which of the moments is the most big, hourly or counterclockwise moment? b) Will the disk be in equilibrium? Why ? c) In which direction will the disk rotate, clockwise or counterclockwise?


1
Expert's answer
2021-02-23T10:06:05-0500

a) The moments (torque) are given as follows:


"\tau_1 = F_1d_1 = 10N\cdot 0.08m = 0.8\space N\cdot m\\
\tau_2 = F_2d_2 = 8N\cdot 0.05m = 0.4\space N\cdot m"

Thus, the first one is the most big.


b) Assuming that these moments are applied in opposite directions, the disk will not be in equilibrium. The nessesary condition for a body to be in equilibrium is the following: sum of all moments applied to it should be zero. Since "\tau_1>\tau_2", this is not the case in this problem.


c) Since "\tau_1>\tau_2", disk will rotate in the direction the force "F_1" pulls it.


Answer. a) "\tau_1", d) No, c) in the direction of force "F_1".